1 mol of N2 and 4 mol of H2 are allowed to react in a vessel and after reaction H2O is added. Aqueous solution required 1 mol of HCl. Mole fraction of H2 in gaseous mixture after reaction is

Grade:11

1 Answers

Tanvi Sehra
5 years ago
N2+3H2→2NH3 Initial. 1. 4. 0 Final. 0.5. 2.5. 1Only one mole is formed as it required one moles of HCl only.NH3+H20 → NH4OH NH4OH + HCl → NH4Cl + H2OGaseous products are H2 and N2 So moles fraction of H2 =2.5÷( 0.5+2.5) = 5/6Ans = 5/6